OpenStudy (anonymous):

linear equation 3x+y=0 2x+y=-3.....so you add 3x+2x and get 5x then you will add y+y and 2 y correct then you have -3-0 which becomes 5x+2y=-3 then I am lost please help

OpenStudy (aripotta):

wait, what's the question?

OpenStudy (anonymous):

lol yeah you are back

OpenStudy (anonymous):

linear equation by graphing 3x+y=0 2x+4=-3

OpenStudy (aripotta):

you wanna change those equations from standard form to slope-intercept form

OpenStudy (anonymous):

dont you add the 3x +2x=5x then the y+y=2y so 5x+2y=-3 right that 4 above should be a y

OpenStudy (aripotta):

um, no. you don't do that lol. you have to change them from standard to slope-intercept form

OpenStudy (aripotta):

then you can graph them easily

OpenStudy (anonymous):

so that is the y2-y1 and the x2-x1 run and rise ordeal

OpenStudy (aripotta):

no. that was the slope formula

OpenStudy (aripotta):

standard form: ax + by = c slope-intercept form: y = mx + b the equations of the lines are currently in standard form, so you have to change them into slope-intercept

OpenStudy (aripotta):

i'll do the first equation for you

OpenStudy (aripotta):